OpenStudy (alivejeremy):

The perimeter of a rectangle is 44 inches. If the width of the rectangle is 7 inches, what is the length? 15 inches 37 inches 22 inches 30 inches

OpenStudy (alivejeremy):

can someone help

OpenStudy (michele_laino):

hint: using your data, we can write this: \(7+x+7+x=44\) where \(x\) is the requested length

OpenStudy (michele_laino):



OpenStudy (anonymous):

:)

OpenStudy (alivejeremy):

why u smile

OpenStudy (anonymous):

the answer is 15 :) and idk I was thinking duh lol

OpenStudy (alivejeremy):

can u help me wit more

OpenStudy (mathmale):

Please post each question separatelly, OK?

OpenStudy (anonymous):

Sure :) more like this? or what but yeah through them at me :)

OpenStudy (anonymous):

you dont have to I will answer u Q here

OpenStudy (alivejeremy):

The perimeter of a rectangle is 38 inches. If the length of the rectangle is 14 inches, which equation could be used to find the width, x? x + 2(14) = 38 2(x + 14) = 38 14(x + 2) = 38 x + 14 = 38

OpenStudy (anonymous):

the second one :)

OpenStudy (michele_laino):

hint: if I call the width with \(x\), then I can write this: \(x+14+x+14=38\)

OpenStudy (alivejeremy):

yea the second one

OpenStudy (anonymous):

any more?

OpenStudy (alivejeremy):

yea

OpenStudy (anonymous):

k what u got?

OpenStudy (alivejeremy):

i never got the anwser for the second one

OpenStudy (anonymous):

its the second one one the second question :) sorry if I confused u but yeah B

OpenStudy (anonymous):

*on*

OpenStudy (alivejeremy):

ok u ready for the next one

OpenStudy (anonymous):

::)

OpenStudy (alivejeremy):

Riley is saving up to buy a new television. She needs a total of $1,067.75. Riley has saved $200.95 already and earns $78.80 per week at her job. How many weeks does Riley need to work to save enough money to buy the new television? 9 12 11 10

OpenStudy (alivejeremy):

?

OpenStudy (anonymous):

sorry :) not ::)

OpenStudy (alivejeremy):

u don't get it

OpenStudy (anonymous):

answer 11

OpenStudy (alivejeremy):

can we do more

OpenStudy (anonymous):

sure! but prob only 1 or 2 more I gtg soon :)

OpenStudy (alivejeremy):

ok

OpenStudy (alivejeremy):

Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 9 to the number and multiplied the sum by 2, it would equal 8. Which equation could Stephanie use to find the number, x? 2(x + 9) = 8 2(x) + 9 = 8 9(x + 2) = 8 2 + (x + 9) = 8

OpenStudy (alivejeremy):

do u get it

OpenStudy (anonymous):

B

OpenStudy (alivejeremy):

yea i think it B to

OpenStudy (alivejeremy):

its*

OpenStudy (anonymous):

k anymore I prob only do 1 more lol

OpenStudy (alivejeremy):

Mary spent a total of $322.43 for a party. She spent $200.71 on food, plus an additional $30.43 for each hour of the party. How long was the party? 4 hours 6 hours 5 hours 3 hours

OpenStudy (alivejeremy):

last one

OpenStudy (anonymous):

k holds on requires some math

OpenStudy (anonymous):

4 hours/ A

OpenStudy (anonymous):

medal?

OpenStudy (alivejeremy):

thx for the help byee

OpenStudy (anonymous):

okay thnx :)